OpenStudy (anonymous):

A cylinder labeled X has a height of 16 inches and a radius of 10 inches. Another cylinder labeled Y has a height of 32 inches and a radius of 20 inches. The volume of cylinder Y is _______ times the volume of cylinder X.

OpenStudy (anonymous):

According to Volume of cylinder: \[V = Pr^{2}h\] So: Vx = P * 100 * 16 = 1600P Vy = P * 400 * 32 = 12800P Vy/Vx = 12800P/1600P = 8 So the volume of cylinder Y is 8 times the volume of cylinder X.

OpenStudy (anonymous):

Thank you!!